OpenStudy (anonymous):

Find the measure of arc DC if measure of arc BC is equal to 51° and angle A measures 21

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (anonymous):

what do you think

OpenStudy (anonymous):

i think its 93

OpenStudy (anonymous):

close

OpenStudy (anonymous):

then it must be 72

OpenStudy (anonymous):

@Michele_Laino 21 = (51-x)/2

OpenStudy (anonymous):

that would be the equation to solve for the measure of DC

OpenStudy (michele_laino):

hint: your problem can be modeled by the subsequent drawing:

OpenStudy (michele_laino):

where O is the center of your circumference

OpenStudy (michele_laino):

now what is the measure od the angle <OQD?

OpenStudy (anonymous):

i havent done angles in a long time. this question came out from the blue

OpenStudy (michele_laino):

hint: angles OQD and AQC are vertical angles

OpenStudy (michele_laino):

we have that the measure of angle OQD is 69 degrees, so the mesure of the angle OQB is 180-69=...?

OpenStudy (anonymous):

111 i meant

OpenStudy (michele_laino):

that's right!

OpenStudy (michele_laino):

so the measure of the angle OBQ, is: 180-(51+111)=...?

OpenStudy (anonymous):

18

OpenStudy (michele_laino):

that's right!

OpenStudy (anonymous):

your right

OpenStudy (michele_laino):

Now also the measure of angle ODQ is 18 degrees, since the triangle BOD, is an isosceles triangle

OpenStudy (michele_laino):

so the requaeste measure x, have to checks this equality: x+18+69=180, where x is the measure of the angle COD. What is x?

OpenStudy (michele_laino):

oops..the requested* measure

OpenStudy (anonymous):

x = 93

OpenStudy (michele_laino):

That's right! sorry for my grammar errors

OpenStudy (michele_laino):

so, what is the right option?

OpenStudy (anonymous):

93 would be correct?

OpenStudy (michele_laino):

yes! It is correct!

OpenStudy (anonymous):

@Michele_Laino last ones

OpenStudy (michele_laino):

I'm trying to solve your exercise, please wait

OpenStudy (michele_laino):

Hint: 2 events are mutually exclusive, when the occurring of one excludes the occurring of the other

OpenStudy (michele_laino):

so, if I draw a king, I can not draw a queen, and vice versa

OpenStudy (michele_laino):

the event king of heart exists, namely it can occur

OpenStudy (michele_laino):

and a club is a black card

OpenStudy (anonymous):

so the answer would be D ?

OpenStudy (michele_laino):

a face card means a Queen, or a King or a Jack, I think

OpenStudy (anonymous):

so all of them are connected somehow

OpenStudy (michele_laino):

I think that the in the first option there are two events mutually exclusive

OpenStudy (michele_laino):

namely, a card can be a King, or can be a Queen

OpenStudy (anonymous):

hmm, im confused. i cant agree on which is the right option

OpenStudy (michele_laino):

why?

OpenStudy (anonymous):

which one would be the correct option! A?

OpenStudy (michele_laino):

yes! I think so!

OpenStudy (michele_laino):

please wait a moment I will be back between 10 minutes

OpenStudy (anonymous):

have a couple minutes left to submit the assaignment. mind helping with the last question. @Michele_Laino . i choose A as the answer. is that correct

OpenStudy (michele_laino):

I'm sorry, since we have to know something of genetics here

OpenStudy (anonymous):

@Michele_Laino its probablity. not anything about gentiics. can you take a look at the chart and it will explain it a little

OpenStudy (michele_laino):

ok! please wait a moment I'm trying to solve your exercise, and then I will explane my solution

OpenStudy (michele_laino):

explain*

OpenStudy (michele_laino):

here is my reasoning: 1) the probability to pick a person of group A is: 264/311= 0.8488 2) the probability to pick a person with a Rh positive is: 264/629= 0.4197 3) the probability to pick a person with group A "AND" Rh positive, is: 264/740=0.3567 So the requested probability is given by the subsequent computation: (0.8488+0.4197)-0.3567=...? I have used this formula P(A union B)=P(A)+P(B)-P(A intersection B) where A, and B are two generic events